# RET-CLIP: A Retinal Image Foundation Model Pre-trained with Clinical Diagnostic Reports

### v1

@@ 1 Introduction
To alleviate the above issues, we have the following contributions in this paper: Firstly, we propose a vision-language foundation model for CFPs, named RET-CLIP, which can achieve state-of-the-art performance across diverse tasks and datasets. Secondly, to our best knowledge, this is the first attempt to leverage original clinical diagnostic reports with minimal modifications to build a foundation model, enriching the model’s visual encoding capabilities with practicality and authenticity. The diagnostic reports in Chinese are also included, extending linguistic versatility of the research domain beyond English. Lastly, inspired by clinical practice, we refine our pre-training process to conduct contrastive learning at monocular and patient levels, confirming the significance of this training strategy.
@@ 2.1 Data Collection and Preprocessing
Dataset acquisition. We collected a dataset of retina fundus binocular images-text triplets (RET-Clinical) at the patient level for RET-CLIP. The dataset includes a total of 193,865 samples from *** Hospital. Each patient’s triplet includes two CFPs for left and right eyes, alongside a clinical diagnostic report. These reports, acknowledged by experienced ophthalmologists, are subjected to thorough data cleansing and manual scrutiny to standardize the text, making it suitable for model training. It’s important to highlight that the preprocessing of text data is limited to basic text standardization, avoiding the need for advanced clinical knowledge or modifications that may alter the original content or meaning.
Data preprocessing and augmentation. For the CFPs, all of them are resized to $512\times 512$. The augmentation includes random crop followed by resizing to $224\times 224$, random horizontal flipping, color jitter, and image normalization. For diagnostic reports, we focused on expanding abbreviations to their full forms, unitizing any mixed-language content to align with our text encoder’s language capabilities, and ensuring the text was coherent and grammatically correct.
@@ 2.2 Model Architecture
Text encoder. For a given patient’s diagnostic report $T$, a BERT-based [4] encoder $\Phi_{t}(\cdot)$ is implemented to encode the clinical descriptions with a text token of length $l$:
$\boldsymbol{t}=\Phi_{t}(\boldsymbol{T})\in\mathbb{R}^{l\times d},\mathrm{~{}}\boldsymbol{t}^{p}=\boldsymbol{t}_{0}\in\mathbb{R}^{d},$ (3)
where $\boldsymbol{t}$ denotes the sentence embedding, $\boldsymbol{t}_{0}$ denotes the embedding for [CLS] token, and $\boldsymbol{t}^{p}$ refers to the patient-level textual information. We then implement three stacked two-layer nonlinear MLPs $F_{l}(\cdot)$, $F_{r}(\cdot)$, $F_{p}(\cdot)$ to decouple $t^{p}$ into textual features representing the left eye, right eye, and patient level, termed as $\boldsymbol{T}^{l}$, $\boldsymbol{T}^{r}$, and $\boldsymbol{T}^{p}$, respectively:
$\boldsymbol{T}^{l}=F_{l}(\boldsymbol{t}^{p}),\mathrm{~{}}\boldsymbol{T}^{r}=F_{r}(\boldsymbol{t}^{p}),\mathrm{~{}}\boldsymbol{T}^{p}=F_{p}(\boldsymbol{t}^{p}),\mathrm{~{}}\boldsymbol{T}^{l},\boldsymbol{T}^{r},\boldsymbol{T}^{p}\in\mathbb{R}^{d}.$ (4)
@@ 3.4 Ablation study
The results, as shown in Table 5, confirm the significance of optimizing objectives at both monocular and patient levels. As previously discussed, the combination of the global information provided at the patient level with the finer-grained features contributed at the monocular level is essential to achieve optimal performance.

### v2

@@ 1 Introduction
To alleviate the above issues, we have the following contributions in this paper: Firstly, we propose a vision-language foundation model for CFPs, named RET-CLIP, which we believe is the first attempt to leverage clinical diagnostic reports to build a retinal foundation model, enriching the model’s visual encoding capabilities with practicality and authenticity. The diagnostic reports in Chinese are included, extending the linguistic versatility of the research domain beyond English. Secondly, a novel strategy is proposed to decouple the information of left and right eyes in diagnostic reports, which is a simple yet effective paradigm for building a retinal foundation model. In practical scenarios, diagnostic reports are usually patient-level, mixing information from both eyes, which brings a big challenge for directly using CLIP to build foundation models. The proposed monocular and patient-level contrastive learning approach can handle this challenge in the ophthalmology domain. Lastly, our model achieves state-of-the-art performance across diverse tasks and datasets, confirming the effectiveness of the proposed training strategy.
@@ 2.1 Data Collection and Preprocessing
Dataset acquisition. We collected a dataset of retina fundus binocular images-text triplets (RET-Clinical) at the patient level for RET-CLIP. The dataset includes a total of 193,865 samples from Beijing Tongren Hospital, Beijing, China. Each patient’s triplet includes two CFPs for left and right eyes, alongside a clinical diagnostic report.
Data preprocessing and augmentation. For the CFPs, all of them are resized to $512\times 512$. The augmentation includes random crop followed by resizing to $224\times 224$, random horizontal flipping, color jitter, and image normalization. For diagnostic reports, we focus on correcting typos and consecutive punctuation errors caused by human input, restoring abbreviations to their full expressions, unifying mixed Chinese and English expressions into Chinese to align with our text encoder’s language capabilities, and ensuring the text is coherent and grammatically correct by manual scrutiny. It’s important to highlight that the preprocessing of text data only involves basic text standardization mentioned above, avoiding the need for advanced clinical knowledge or modifications that may alter the original content or meaning.
@@ 2.2 Model Architecture
Text encoder. For a given patient’s diagnostic report $\mathcal{T}$, a BERT-based [4] encoder $\Phi_{t}(\cdot)$ is implemented to encode the clinical descriptions with a text token of length $l$:
$\boldsymbol{T}=\Phi_{t}(\mathcal{T})\in\mathbb{R}^{l\times d},\mathrm{~{}}\boldsymbol{T}_{0}\in\mathbb{R}^{d},$ (3)
where $\boldsymbol{T}$ denotes the sentence embedding, $\boldsymbol{T}_{0}$ denotes the embedding for [CLS] token. We then implement three stacked two-layer nonlinear MLPs $F_{l}(\cdot)$, $F_{r}(\cdot)$, $F_{p}(\cdot)$ to decouple $\boldsymbol{T}_{0}$ into textual features representing the left eye, right eye, and patient level, termed as $\boldsymbol{T}^{l}$, $\boldsymbol{T}^{r}$, and $\boldsymbol{T}^{p}$, respectively:
$\boldsymbol{T}^{l}=F_{l}(\boldsymbol{T}_{0}),\mathrm{~{}}\boldsymbol{T}^{r}=F_{r}(\boldsymbol{T}_{0}),\mathrm{~{}}\boldsymbol{T}^{p}=F_{p}(\boldsymbol{T}_{0}),\mathrm{~{}}\boldsymbol{T}^{l},\boldsymbol{T}^{r},\boldsymbol{T}^{p}\in\mathbb{R}^{d}.$ (4)
@@ 3.1 Tasks and Datasets
For the IDRIR, the entire dataset is officially divided into a test set comprising 20% of the data, with the remaining 80% designated as the training set. In our experiments, we further split the training set into a training set and a validation set using a 4:1 ratio. Similarly, for the PAPLA, we follow the official partitioning method, which aligns with the approach described above. Regarding the RFMID, the official division includes distinct sets for training, validation, and testing; we adhere to this official partitioning. For all other datasets, we divide them into training, validation, and test sets using a 0.56:0.14:0.3 ratio, following RETFound’s [25] partitioning method. For all datasets, samples within each category are initially distributed based on the specified proportions before being combined to ensure consistent category distribution across the training, validation, and test sets.
@@ 3.4 Ablation study
The results, as shown in Table 5, confirm the effectiveness of optimizing objectives at both monocular and patient levels. As previously discussed, the combination of the global information provided at the patient level with the finer-grained features contributed at the monocular level is essential to achieve optimal performance.
@@ 4 Conclusion
{credits}
4.0.1 Acknowledgements
This research work is supported by Beijing Natural Science Foundation (Grant No. IS23112), National Natural Science Foundation of China (NSFC) (GrantNo. 82072007), and Beijing Institute of Technology Research Fund Program for Young Scholars.
[…]
The authors have no competing interests to declare that are relevant to the content of this article.